ALGAE

Algae Graphical User Interface

Peter Eisenmann peisen@cosy.sbg.ac.at

Clemens Bauer cbauer@cosy.sbg.ac.at Thomas Frommwald tfromm@cosy.sbg.ac.at

Daniel Kadir dkadir@cosy.sbg.ac.at Michael Liedlgruber mliedl@cosy.sbg.ac.at

Head of the course: Roland Schwaiger

May 15, 2000

Chapter 1

GUI Description

1.1 Configuration Window (n/a yet)

The program starts with the configuration window where all parameters relevant for the simulation are adjusted. There is a dropdown list where one of several preset configurations may be selected. If a specific configuration is input, it can be saved as a new preset.

1.2 Main GUI

1.2.1 Aquarium

This is the world where our simulation takes place. It shows the agents, pheromon tracks and substrat parts. The current light intensity is shown by an icon which changes color (n/a yet)

When the sim is paused, it is easy to select an specific agent; a new agent information window is opened then (see below)

1.2.2 MenuBar

FileMenu

• New (starts a new simulation and resets the parameters)

• Load configuration (what the name says)

• Save configuration (what the name says)

• Save sim status (saves everything concerning the world and its agents)

• Load sim status (opens a saved status)

• Quit (goodbye)

WindowsMenu • List of all open windows

• Show simulation info window (see below) HelpMenu

• Help contents

• About

1.2.3 ToolBar

We control most parts of the simulation through the toolbar. It is possible to start, pause and restart the sim and to advance step by step. We can raise or lower the zoom factor, from 1 pixel for each algae to a 10*10 pixel view. The speed of the simulation is controlled via the delay slider; one can also turn off the graphic output completely with the display switch. The second slider controls the intensity of the light source. To simulate the day/night cycle, this slider can be disabled with the day/night cycle switch. Display of pheromon and substrat can be turned on or off with two buttons (n/a yet)

1.2.4 StatusBar

By default, it shows the current number of agents and the elapsed time. When a specific point in our world is selected, we get the following information:

• Is there substrat at this place ?

• Are any phermomon tracks here ?

• Light intensity at this place ?

• The XY Position ?

1.3 Agent Information Window (n/a yet)

• Current sensor values

• The neural network

• The agent’s genotype (bitstring)

• The speed, direction and action (i.e. eating) of the algae

• The energy level, age and current position

• Other information ?

1.4 Simulation Information Window (n/a yet)

Here we get several statistics about the status of our population and the simulation itself.